### v1

@@ Appendix A Full MMS by Removing $n-2$ Bads for Chore Allocation
For any fair division instance of chores with additive costs, at least $m-n+2$ items can be allocated while maintaining a full MMS allocation.
@@
As in the proof of Lemma 4, if all agents are allocated by a perfect matching in $G$, and then we don’t need to throw away any items. If we match a subset of the agents by bundles in $G$, then the remaining agents do not have any edges to the allocated bundles, which means that if $k$ agents were allocated, the rest of the agents have a cost of at most $|N|-k$ for the remaining bundles. Thus, we can apply Lemma 22 and get that at most $|N|-k-1$ items are thrown away. Since $k\geq 1$, we get that the number of allocated items is at least $|M|-|N|+2$. ∎
@@ Appendix D MMS with copies for $k$-demand valuations
This suggests a variant of algorithm $\mathsf{BagFill\mbox{-}with\mbox{-}Copies}$ for $k$-demand valuations. Upon processing a new good $g$, and adding it to $B$, don’t assign the entire bag $B$ to an agent with $v(B)\geq 1$. Instead, choose some set $S\subseteq B$, $|S|=k$ with $v(S)\geq 1$. Some such set $S$ must exist (as $v(B)\geq 1$ and $v$ is $k$-demand). Moreover, the last good added to the bag $g$ must be in $S$, otherwise the bag $B^{\prime}=B\setminus\{g\}\subset B$, considered just prior to the arrival of $g$, would also have had value $\geq 1$. Give one of the agents with $v(B)\geq 1$ some $k$-subset of $B$, $S\subset B$, for which $v(S)\geq 1$. Return all goods in $B\setminus S$ to the pool of goods yet to be processed, and assign $B=\{g\}$, the last good added to that triggered this allocation, this creates a copy of $g$. Also, such duplications can be done at most $n-1$ times and no good is duplicated more than once.

### v2

@@ Appendix A Full MMS by Removing $n-2$ Bads for Chore Allocation
For any fair division instance of chores $\mathcal{I}=(N,M,\{v_{i}\}_{i\in N})$ with additive costs, Algorithm $\mathsf{Match\mbox{-}n\mbox{-}Fill\mbox{-}for\mbox{-}Chores}(\mathcal{I})$ outputs a full MMS allocation and allocates at least $|M|-|N|+2$ items.
@@
As in the proof of Lemma 4, if all agents are allocated by a perfect matching in $G$, and then we do not need to throw away any items. If we match a subset of the agents by bundles in $G$, then the remaining agents do not have any edges to the allocated bundles, which means that if $k$ agents were allocated, the rest of the agents have a cost of at most $|N|-k$ for the remaining bundles. Thus, we can apply Lemma 23 and get that at most $|N|-k-1$ items are thrown away. Since $k\geq 1$, we get that the number of allocated items is at least $|M|-|N|+2$. ∎
@@ Appendix D MMS with copies for $k$-demand valuations
This suggests a variant of algorithm $\mathsf{BagFill\mbox{-}with\mbox{-}Copies}$ for $k$-demand valuations. Upon processing a new good $g$, and adding it to $B$, do not assign the entire bag $B$ to an agent with $v(B)\geq 1$. Instead, choose some set $S\subseteq B$, $|S|=k$ with $v(S)\geq 1$. Some such set $S$ must exist (as $v(B)\geq 1$ and $v$ is $k$-demand). Moreover, the last good added to the bag $g$ must be in $S$, otherwise the bag $B^{\prime}=B\setminus\{g\}\subset B$, considered just prior to the arrival of $g$, would also have had value $\geq 1$. Give one of the agents with $v(B)\geq 1$ some $k$-subset of $B$, $S\subset B$, for which $v(S)\geq 1$. Return all goods in $B\setminus S$ to the pool of goods yet to be processed, and assign $B=\{g\}$, the last good added to that triggered this allocation, this creates a copy of $g$. Also, such duplications can be done at most $n-1$ times and no good is duplicated more than once.
@@ Appendix E Example for Algorithm $\mathsf{BagFill\mbox{-}with\mbox{-}Copies}$
Appendix F Picking Sequence Fails for Allocation with Copies
Consider three agent $\{a,b,c\}$ with additive valuations over three goods $x,y,z$. Let the ordinal valuations be as follows:
$v_{a}:$ $x\succ z\succ y$
$v_{b}:$ $x\succ z\succ y$
$v_{c}:$ $x\succ y\succ z$
Consider the ordered counterpart (see Definition 4) of this instance and the allocation $A^{\prime}$ below, in which every good is copied once.
$A^{\prime}_{a}=\{1,2\},\qquad A^{\prime}_{b}=\{1,3\},\qquad A^{\prime}_{c}=\{2,3\}$
We interpret this (virtual) allocation as inducing the following value guarantee: for each good $j$ that agent $i$ receives in $A^{\prime}$, she is entitled to a good whose value is at least her $j$ most-valuable good in the original instance. E.g., agent $b$ in the above must receive $x$ and either of $y$ or $z$.
One can generalize the picking sequence algorithm ([16], see also Proposition 2), to a setting with copies in two ways. Assume we have a virtual allocation $A^{\prime}$, where each good has at most $k$ extra copies.
One way is to run $k+1$ instances of the picking sequence algorithm, where in each run each good in $M$ is allocated at most once. The first agent to pick is an agent which holds (one of the instances of) the virtual good $1$, the second holds $2$ and so on. Another way, is to allow all agents which hold good $1$ to pick first, then all agents which hold good $2$, and so on. In this interpretation agent $i$ picks the most-valued good $g$ such that $g\notin A_{i}$, and $g$ was allocated no more than $k+1$ times (so each original good is copied at most $k$ times as well). Below we show that neither of these interpretations work if we break the ties arbitrarily. In fact, the first one fails for any tie-breaking rule, see Example 2.
The failure of the two generalizations follows from the fact that an agent is not allowed to have two instances of the same good. If valuations were such that any $q$ copies of any good $g$ would yield a value of $q\cdot v(g)$, the same proof as in [16] would carry over to this setting. We show examples in which the two generalizations fail when at most one extra copy is made (i.e., for $k=1$). For the first generalization, the example below shows that breaking ties arbitrarily may fail.
Example 1.
Consider the valuations and (virtual) allocation presented at the beginning of the section. Assume ties are broken such that the picking sequences in the two rounds are $a,c,b$ and $b,a,c$, respectively. The first run results in $(A_{a},A_{b},A_{c})=(\{x\},\{z\},\{y\})$. In the second run, since the picking sequence is $b,a,c$, before agent $c$ gets to pick the allocation would be $(A_{a},A_{b},A_{c})=(\{x,z\},\{z,x\},\{y\})$ which leaves no item for $c$ to pick in the last round.
For the second generalization, consider the following example, in which the value guarantees do not hold, regardless of tie-breaking.
Example 2.
$v_{a}:$ $x\succ y\succ z$
$v_{b}:$ $z\succ y\succ x$
$v_{c}:$ $z\succ y\succ x$
with the allocation
$A^{\prime}_{a}=\{1,2,3\},\qquad A^{\prime}_{b}=\{1\},\qquad A^{\prime}_{c}=\{2,3\}$
Agents which hold good 1 pick: $(A_{a},A_{b},A_{c})=(\{x\},\{z\},\emptyset)$.
Agents which hold 2 pick: $(A_{a},A_{b},A_{c})=(\{x,y\},\{z\},\{z\})$.
Agents which hold 3 pick: Agent a cannot pick another good.